# Recall, Robustness, and Lexicographic Evaluation

### v5

@@ Appendix A Metric Properties
A.1. Properties of Top-Heavy Recall-Level Metrics
@@
See the first case in the proof Theorem A.1. ∎
Theorem A.3 ().
@@
See the second case in the proof Theorem A.1. Moreover, if $z(i,m)$ is strictly positive, then Equation 13 will always be strictly positive. ∎
Theorem A.4 ().
@@
A.2. Properties of Total Search Efficiency
Let $\mu$ be TSE with an arbitrary exposure function $e$ and normalization function $z$. Because TSE is a top-heavy recall-level metric, we know that it satisfies all of the properties in Section A.1 except those conditional on $z(i,m)>0$, since $z_{\text{SL}_{3}}(i,m)=0$ when $i<m$.
A.3. Properties of lexirecall
Theorem A.7 ().
@@ A.3. Properties of lexirecall
See the second case in the proof Theorem A.7. ∎
Theorem A.10 ().
@@ Appendix B Robustness
B.1. Worst-case and Total Search Efficiency
@@
B.2. Leximin and Lexicographic Recall
@@
(a) Users. Partitioning $\mathcal{W}$ when $k=3$. $\mathcal{W}$ is partitioned into $\mathcal{W}_{\geq k}-\mathcal{W}_{>k}$ (double circle), $\mathcal{W}_{>k}$ (dashed circle), and $\mathcal{W}-\mathcal{W}_{\geq k}$ (colored circles, described in the proof of Lemma B.3). This figure best rendered in color.
@@
Theorem B.1 means that $\varrho_{m^{+}}=q_{m}$ and $\varrho^{\prime}_{m^{+}}=q^{\prime}_{m}$ and the implication is true.
@@
Because $q\succ_{\text{leximin}}q^{\prime}$, we know that there exists a $k$ such that $q_{k}>q^{\prime}_{k}$ and $\forall i\in[k+1\mathrel{{.}\,{.}}\nobreak m],q_{i}=q^{\prime}_{i}$ (as depicted in Figure 18). We know from Lemma B.3 that the lowest-ranked, non-tied users are in $\mathcal{W}_{\geq k}-\mathcal{W}_{>k}$.
@@
We know from the proof of Theorem B.2, the lowest ranked items in $\phi$ and $\phi^{\prime}$ correspond to $e(p_{m})=q_{m}$ and $e(p^{\prime}_{m})=q^{\prime}_{m}$, respectively, and so the proof holds in this case.
@@
Appendix C Optimal Rankings
@@
Appendix D Number of Ties
Theorem D.1 ().
@@ Appendix D Number of Ties
Theorem D.3 ().
[…]
@@
Theorem D.4 ().
[…]
@@
The proof follows that of Theorem D.3, substituting $m$ for $k$. ∎

### v6

@@ Appendix B Metric Properties
B.1. Properties of Top-Heavy Recall-Level Metrics
@@
See the first case in the proof Theorem B.1. ∎
Theorem B.3.
@@
See the second case in the proof Theorem B.1. Moreover, if $z(i,m)$ is strictly positive, then Equation 13 will always be strictly positive. ∎
Theorem B.4.
@@
B.2. Properties of Total Search Efficiency
Let $\mu$ be TSE with an arbitrary exposure function $e$ and normalization function $z$. Because TSE is a top-heavy recall-level metric, we know that it satisfies all of the properties in Section B.1 except those conditional on $z(i,m)>0$, since $z_{\text{SL}_{3}}(i,m)=0$ when $i<m$.
B.3. Properties of lexirecall
Theorem B.7.
@@ B.3. Properties of lexirecall
See the second case in the proof Theorem B.7. ∎
Theorem B.10.
@@ Appendix C Robustness
C.1. Worst-case and Total Search Efficiency
@@
C.2. Leximin and Lexicographic Recall
@@
(a) Users. Partitioning $\mathcal{W}$ when $k=3$. $\mathcal{W}$ is partitioned into $\mathcal{W}_{\geq k}-\mathcal{W}_{>k}$ (double circle), $\mathcal{W}_{>k}$ (dashed circle), and $\mathcal{W}-\mathcal{W}_{\geq k}$ (colored circles, described in the proof of Lemma C.3). This figure best rendered in color.
@@
Theorem C.1 means that $\varrho_{m^{+}}=q_{m}$ and $\varrho^{\prime}_{m^{+}}=q^{\prime}_{m}$ and the implication is true.
@@
Because $q\succ_{\text{leximin}}q^{\prime}$, we know that there exists a $k$ such that $q_{k}>q^{\prime}_{k}$ and $\forall i\in[k+1\mathrel{{.}\,{.}}\nobreak m],q_{i}=q^{\prime}_{i}$ (as depicted in Figure 19). We know from Lemma C.3 that the lowest-ranked, non-tied users are in $\mathcal{W}_{\geq k}-\mathcal{W}_{>k}$.
@@
We know from the proof of Theorem C.2, the lowest ranked items in $\phi$ and $\phi^{\prime}$ correspond to $e(p_{m})=q_{m}$ and $e(p^{\prime}_{m})=q^{\prime}_{m}$, respectively, and so the proof holds in this case.
@@
Appendix D Optimal Rankings
@@
Appendix E Number of Ties
Theorem E.1.
@@ Appendix E Number of Ties
Theorem E.3.
$\text{\emph{Pr}}(\pi=_{\text{R}_{k}}\pi^{\prime})$ =($={n\choose m}^{-2}\sum_{i=0}^{m}{{k\choose i}^{2}{n-k\choose m-i}^{2}}$
@@
Theorem E.4.
$\text{\emph{Pr}}(\pi=_{\text{RP}}\pi^{\prime})$ =($={n\choose m}^{-2}\sum_{i=0}^{m}{{m\choose i}^{2}{n-m\choose m-i}^{2}}$
@@
The proof follows that of Theorem E.3, substituting $m$ for $k$. ∎